47 - (2x + 5) less than or equal too 3(x+2)+ x
use set-builder notation to describe the complete solution

Answer by mananth(16949)

47 - (2x + 5) <= 3(x+2)+ x
47-2x-5<= 3x+6+x
42-2x<=4x+6
Add 2x to both side
42-2x+2x<=4x+6+2x
42<=6x+6
add -6 to both sides
42-6<=6x+6-6
36<=6x
divide by6 both sides
36/6 <=x
6<=x
{x|x>=6}